# Real-time Tracking in a Status Update System with an Imperfect Feedback Channel

### v1

@@ IV Numerical Results
In this section, we evaluate the performance of the proposed policies: 1) the RVIA-based policy presented in Algorithm 1 and 2) the low complexity policy. In the simulation results, we use the following distortion function: ${g\big{(}X(t),\hat{X}(t)\big{)}=|X(t)-\hat{X}(t)|}$.
In the following, we first study the effect of truncation on the performance of the RVIA-based policy, then we study the structure of the proposed policies, and finally, we analyze the performance of the proposed policies with respect to different system parameters.
IV-A RVIA-based Policy
In Fig. 2, we evaluate the effect of belief-state truncation on the performance of the RVIA-based policy (presented in Algorithm 1) with $\epsilon=1e-4$. Fig. 2 depicts the average cost versus the number of consecutive transmission attempts without receiving ACK, i.e., $m$, for different values of probabilities of successful reception in the forward and backward channels. As can be seen, when the channels’ condition is bad, e.g., $p_{s}=p_{f}=0.2$, the performance of the derived policy converges slowly with $m$. This is because, with a small value of $p_{s}$ and $p_{f}$, the number of consecutive transmission attempts without receiving ACK increases, and thus the truncated belief space needs to include states associated with large values of $m$ to be a good approximation for the actual belief space. However, when the probability of receiving ACK increases, the average cost $\bar{C}$ converges quickly with $m$. Fig. 2 further shows that increasing $m$ beyond the value $m=4$ leads to negligible improvement in the objective function. Hence, a value of $m=4$ is a reasonable choice to truncate the belief-state space for the considered setting.
[…]
Figure 2: The average cost $\bar{C}$ versus $m$ for different reliability’s of the forward and feedback channels with $p=0.7$, $\mu=0.5$, $B=3$, and $N=3$.
Fig. 3 illustrates the evolution of the average cost $\bar{C}$ as a function of time for different probabilities of successful reception $p_{s}$. From Fig. 3, we see that the average cost decreases when $p_{s}$ increases. This is due to the fact that by improving the channel condition the number of successful receptions increases, and consequently, the number of slots at which the system is in an erroneous state decreases.
Figure 3: The (instantaneous) average cost $\bar{C}$ for different values of successful reception in the forward channel for the RVIA-based policy with respect to time slots where $p=0.7$, $p_{f}=0.7$ $\mu=0.5$ $B=3$, and $N=3$.
@@ IV-B The Structure of Proposed Policies
Fig. 4 visualizes the structure of the RVIA-based policy for different values of the energy arrival rates $\mu$ where the source is binary, and the source state is $X=1$. The plot is drawn with respect to the battery’s energy level $b$, and the belief associated with $\hat{X}=1$, i.e., $\rho_{1}$. As can be seen, the policy has a threshold structure with respect to $b$ and $\rho_{1}$. For example, when $\mu=0.2$ and $\rho_{1}=0.5598$, the action decision for $b\geq 3$ is transmission $a=1$ and for $b<3$ is staying idle $a=0$. In addition, Fig. 4 shows that by increasing $\mu$, the number of belief states that the policy decides to transmit increases. This is because increasing $\mu$ results in having more energy in the system.
@@
Figure 4: Structure of the RVIA-based policy for a binary source with respect to the belief and energy level of the battery $b$ for three different energy arrival rate $\mu$, where $X=1$, $p_{s}=0.6$, $p=0.7$, $p_{f}=0.7$, $B=3$, and $N=2$.
@@
Fig. 5 visualizes the structure of the RVIA-based policy for different source self-transition probabilities $p$, where the source is binary, and the source state is $X=1$. The plot is drawn with respect to $b$ and $\rho_{1}$. The figure shows that when $p$ is small, e.g., $p=0.5$, the policy tends not to transmit. This is because there is a high chance that the source would change its state at the next slot, and the currently transmitted status update cannot decrease the average distortion at the next slot. However, when $p$ increases, the policy tends to have transition attempts more frequently. It can also be seen that the policy has a threshold structure with respect to $b$ and $\rho_{1}$. For instance, when $p=0.9$ and $\rho_{1}=0.8594$, the threshold value for the battery’s energy level is $b=2$. Moreover, the results show that when $p=q$ (e.g., $p=0.5$ and $q=0.5$), in line with [38], the optimal decision is to stay idle.
@@
Figure 5: Structure of the RVIA-based policy for a binary source with respect to the belief and energy level of the battery $b$ for three different source’s self-transition probability $p$, where $X=1$, $p_{s}=0.4$, $\mu=0.7$, $p_{f}=0.5$, $B=3$, and $N=2$.
Fig. 6 demonstrates the structure of decision-making by the low-complexity policy for a given state ${l=\{X=1,b\geq 1,\rho=[0,1,0],d_{-1}=1,X_{-1}=2\}}$ versus the source’s self-transition probability $p$, and the probability of successful reception in the forward channel $p_{s}$. As can be seen, the decision is independent of $p_{s}$ and changes with $p$. For small values of $p$, the policy keeps the transmitter idle, while for a larger value of $p$, it causes the transmitter to send updates most of the time. This is because smaller values of $p$ result in more frequent state changes at the source, increasing the likelihood that discrepancies between the source and the sink are resolved automatically rather than through status updating.
@@
Figure 6: Structure of the low-complexity policy with respect to the source’s self-transition probability $p$, and the probability of successful reception in forward channel $p_{s}$ for belief-state ${l=(X=1,b\geq 1,\rho=[0,1,0],d_{-1}=1,X_{-1}=2)}$ where $B=3$, ${N=3}$.
@@ IV-C Performance Comparison of the Proposed Policies
In this subsection, we compare the proposed policies against each other and a baseline policy. According to the baseline policy, the transmitter sends an update whenever the battery’s energy level is not zero, and it is not certain that the estimation error is zero.
@@
Fig. 7 depicts the average cost $\bar{C}$ of different policies versus the source’s self-transition probability $p$. As can be seen for small values of $p$, e.g., $p=0.4$ or $p=0.5$, the low-complexity policy and the RVIA-based policy perform almost the same, and both outperform the baseline policy. This is because, for small values of $p$, the RVIA-based and low-complexity policies transmit less frequently, while the baseline policy sends updates more frequently (it sends updates whenever the battery’s energy level is not zero, and it is not certain that the estimation error is zero). Actually, for a high-dynamic source, there is a high chance that the source will change its state in the next slot, and the currently transmitted status update cannot decrease the average cost at the next slot. For large values of $p$ (e.g., $p=0.7,p=0.8,p=0.9$), the average cost of the baseline policy is very close to the RVIA-based policy. This is because since the source state changes slowly for large values of $p$, the energy arrival rate is quite high, and the channel condition is good, making suboptimal decisions does not degrade the baseline policy performance.
[…]
@@
Figure 7: The average cost $\bar{C}$ for the proposed policies with respect to $p$ where $p_{s}=0.6$, $\mu=0.7$, $p_{f}=0.6$, $B=3$, $N=3$.
@@
Fig. 8 shows the average cost $\bar{C}$ versus the energy arrival rate $\mu$ for different policies. As it can be seen, for small values of $\mu$, the RVIA-based policy outperforms other policies, while for large values of $\mu$, the performance of the policies are very close. This is because in cases where the energy arrival rates $\mu$ are small, there are fewer transmission opportunities, and the low-complexity and baseline policies utilize these opportunities less effectively than the RVIA-based policy.
[…]
@@
Figure 8: The average cost $\bar{C}$ for the proposed policies with respect to the energy arrival rate $\mu$, where $p_{s}=0.4$, $p=0.7$, $p_{f}=0.6$, $B=3$, $N=3$.
@@
Fig. 9 demonstrates the average cost versus the number of source’s states $N$ under different policies. From the figure, the average cost increases with $N$, as expected. Moreover, the RVIA-based policy outperforms other policies. This is because increasing the number of states, in general, leads to a larger estimation error. Consequently, the system needs to use the available energy effectively by considering the state of the system in future time slots. However, the low-complexity policy considers only one slot, and the baseline policy transmits opportunistically.
[…]
@@
Figure 9: The average cost $\bar{C}$ for the proposed policies with respect to numbers of source’s state $N$, where $p_{s}=0.6$, $\mu=0.2$, $p_{f}=0.2$, $p=0.7$, $B=3$, and $N=3$.
@@
In Fig. 10, we examine the impact of the probability of successful reception of ACK $p_{f}$ for different policies. As expected, the average cost decreases by an increase in $p_{f}$. This is because the number of slots that the sink is observable by the transmitter increases, which in turn reduces the uncertainty in the system, leading to an effective decision-making. For example, the performance of the policies is very close for $p_{f}=1$. This is due to the fact that the baseline and low complexity policies can observe the sink and send an update only if there is a mismatch between the source and sink, which prevents wasting energy. This figure shows the effectiveness of the derived RVIA-based policy when the feedback channel condition is bad and its applicability to low-resource wireless networks, where having high-reliable communication is more costly.
[…]
@@
Figure 10: The average cost $\bar{C}$, for the proposed policies with respect to the probability of successful reception of ACK $p_{f}$, where $p_{s}=0.4$, $\mu=0.5$, $p=0.7$, $B=3$, and $N=3$.
@@
Finally, Fig. 11 depicts the average cost with respect to the battery’s capacity for different energy arrival rates. Here, without loss of generality, we use the RVIA-based policy. As expected, the average cost decreases by an increase in $B$. This is because a larger capacity of the battery allows the transmitter to store more energy, compensating for periods of inconsistent energy reception and ensuring continuous operation. The influence of the battery’s capacity on performance diminishes with increasing energy arrival rate. This is because, at a higher energy arrival rate, the transmitter receives energy more frequently. Therefore, the system can rely more on frequent energy arrival rather than large battery capacities.
[…]
@@
Figure 11: The average cost $\bar{C}$ for different values of the battery’s energy arrival $\mu$ with respect to the battery’s capacity $B$, where $p_{s}=0.6$, ${p_{f}=0.7}$, $p=0.7$, and $N=3$.
@@
In Table I, we compare the computational complexity of the proposed policies. The computational complexity of the RVIA is at most $\mathcal{O}(|\mathcal{A}_{l}||\mathcal{L}|^{2})$ [9], where $|\mathcal{A}_{l}|$ is the size of action space and $|\mathcal{L}|$ is the size of belief-state space. The size of the action space is $2$. If we start from a reset belief, after one transmission attempt without receiving ACK, i.e., $m=1$, there are $N-1$ possible beliefs, and for $m=2$, there are $N$ possible beliefs. This procedure can start from any reset belief. Thus, the number of beliefs in the belief space is ${N(N-1)N^{m-1}=(N-1)N^{m}}$. Considering the other elements of the belief-state, the size of the belief-state space becomes $2B(N-1)N^{m+2}$. Therefore, the computational complexity for the offline phase of the RVIA-based policy is $\mathcal{O}([B(N-1)N^{m+2}]^{2})$. In the online phase, the policy needs to find the action from a state-action table, thus the computational complexity is ${\mathcal{O}(1)}$.
The low-complexity policy does not have an offline phase, and in the online phase, it needs to compare the cost associated with two actions, thus the computational complexity for the low-complexity policy in the online phase becomes ${\mathcal{O}(2)}$.
@@
RVIA-based $\mathcal{O}([B(N-1)N^{m+2}]^{2})$ $\mathcal{O}(1)$
Low-complexity - $\mathcal{O}(2)$

### v2

@@ IV Numerical Results
In this section, we evaluate the performance of the proposed policies: 1) the POMDP-based policy presented in Algorithm 1 and 2) the LC policies presented in Section III-D. In the simulation results, we use the following distortion function: ${d\big{(}X(t),\hat{X}(t)\big{)}=|X(t)-\hat{X}(t)|}$.
In the following, we first study the effect of truncation on the performance of the POMDP-based policy, then we study the structure of the proposed policies, and finally, we analyze the performance of the proposed policies with respect to different system parameters.
IV-A POMDP-based Policy
In Fig. 2, we evaluate the effect of belief-state truncation on the performance of the POMDP-based policy (presented in Algorithm 1) with $\epsilon=1e-4$. Fig. 2 depicts the average cost versus the number of consecutive transmission attempts without receiving ACK, i.e., $m$, for different values of probabilities of successful reception in the forward and backward channels. As can be seen, when the channels’ condition is bad, e.g., $p_{s}=p_{f}=0.2$, the performance of the derived policy converges slowly with $m$. This is because, with a small value of $p_{s}$ and $p_{f}$, the number of consecutive transmission attempts without receiving ACK increases, and thus, the truncated belief space needs to include states associated with large values of $m$ to be a good approximation for the actual belief space. However, when the probability of receiving ACK increases, the average cost $\bar{C}$ converges quickly with $m$. Fig. 2 further shows that increasing $m$ beyond the value $m=6$ leads to negligible improvement in the objective function. Hence, a value of $m=6$ is a reasonable choice to truncate the belief-state space for the considered setting.
11112222333344445555666677770.60.60.60.60.70.70.70.7Number of slots without receiving ACK, $m$Average cost, $\bar{C}$$p_{s}=0.2$ $p_{f}=0.2$$p_{s}=0.4$ $p_{f}=0.4$$p_{s}=0.6$ $p_{f}=0.6$
Figure 2: The average cost $\bar{C}$ versus $m$ for different reliabilities of the forward and feedback channels where the source is characterized as $p=0.7$, $N=3$ and the EH module is characterized as $\mu=0.5$, $B=3$.
Fig. 3 illustrates the evolution of the average cost $\bar{C}$ as a function of time for different probabilities of successful reception $p_{s}$ under POMDP-based policy. From Fig. 3, we see that the average cost decreases when $p_{s}$ increases. This is due to the fact that by improving the channel condition the number of successful receptions increases, and consequently, the number of slots at which the system is in an erroneous state decreases.
Figure 3: The (instantaneous) average cost $\bar{C}$ for different values of successful reception in the forward channel for the POMDP-based policy with respect to time slots where $p=0.7$, $N=3$, $\mu=0.5$, $B=3$, and $p_{f}=0.7$.
@@ IV-B The Structure of Proposed Policies
Fig. 4 visualizes the structure of the POMDP-based policy for different values of the energy arrival rates $\mu$ where the source is binary, and the source state is $X=1$. The plot is drawn with respect to the battery’s energy level $b$, and the belief associated with $\hat{X}=1$, i.e., $\rho_{1}$. As can be seen, the policy has a threshold structure [^4] A threshold structure in a policy refers to a decision rule where different actions are taken before and after a certain threshold. with respect to $b$ and $\rho_{1}$. For example, when $\mu=0.2$ and $\rho_{1}=0.5598$, the action decision for $b\geq 3$ is transmission $a=1$ and for $b<3$ is staying idle $a=0$. In addition, Fig. 4 shows that by increasing $\mu$, the number of belief states that the policy decides to transmit increases. This is because increasing $\mu$ results in having more energy in the system.
@@
Figure 4: Structure of the POMDP-based policy for a binary source with respect to the belief and energy level of the battery $b$ for three different energy arrival rate $\mu$, where $X=1$, $p=0.7$, $N=2$, channel reliabilities are defined as $p_{s}=0.6$, $p_{f}=0.7$, and $B=3$.
@@
Fig. 5 visualizes the structure of the POMDP-based policy for different source self-transition probabilities $p$, where the source is binary, and the source state is $X=1$. The plot is drawn with respect to $b$ and $\rho_{1}$. The figure shows that when $p$ is small, e.g., $p=0.5$, the policy tends not to transmit. This is because there is a high chance that the source would change its state at the next slot, and the currently transmitted status update cannot decrease the average distortion at the next slot. However, when $p$ increases, the policy tends to have transition attempts more frequently. It can also be seen that the policy has a threshold structure with respect to $b$ and $\rho_{1}$. For instance, when $p=0.9$ and $\rho_{1}=0.8594$, the threshold value for the battery’s energy level is $b=2$. Moreover, the results show that when $p=q$ (e.g., $p=0.5$ and $q=0.5$), in line with [13], the optimal decision is to stay idle.
@@
Figure 5: Structure of the POMDP-based policy for a binary source with respect to the belief and energy level of the battery $b$ for three different source’s self-transition probability $p$, where $X=1$, $N=2$, $p_{s}=0.4$, $p_{f}=0.5$, $\mu=0.7$, and $B=3$.
Fig. 6 demonstrates the structure of the LC policies for different source self-transition probabilities $p$, where the source is binary, and the source state is $X=1$. The plot is drawn with respect to $b$ and $\rho_{1}$. Fig. 6(a) shows the structure of the energy-agnostic LC policy, and Fig. 6(b) shows the structure of the energy-aware LC policy. As can be seen, the decision is independent of $b$ and changes with $p$ for both policies. For small values of $p$, the decision determined by the policies is to stay idle, while for a larger value of $p$, the decision is to send updates, which is the same as the decision determined by the POMDP-based policy (see Fig. 5). For $p>0.5$, the energy-agnostic LC policy sends updates when there is energy in the battery unless the expected distortion cost is zero, i.e., $\rho_{1}=1$. However, the energy-aware LC policy stays idle when the expected distortion cost is small (e.g., $\rho_{1}\geq 0.7$). This is because the distortion cost is small, sending new updates does not reduce the cost significantly.
(a) Energy-agnostic LC policy
(b) Energy-aware LC policy
@@
Figure 6: Structure of the LC policies for a binary source with respect to the belief and energy level of the battery $b$ for three different source’s self-transition probability $p$, where $X=1$, $N=2$, $p_{s}=0.6$, $p_{f}=0.5$, $\mu=0.7$, and $B=3$.
@@ IV-C Performance Comparison of the Proposed Policies
In this subsection, we compare the proposed policies, i.e., the POMDP-based policy presented in Section III-C and LC policies presented in Section III-D, against each other and two baseline policies. According to the first baseline, the transmitter sends an update whenever the battery’s energy level is not zero, referred to as battery-only (BO) policy. In the second baseline, the transmitter sends an update whenever the battery’s energy level is not zero, and the expected cost in (12) is not zero, referred to as battery-only with redundancy check (BO-RC) policy.
@@
Fig. 7 depicts the average cost $\bar{C}$ of different policies versus the source’s self-transition probability $p$. As it can be seen for small values of $p$, e.g., $p=0.4$ or $p=0.5$, the LC policies and the POMDP-based policy perform almost the same, and outperform the baseline policies. This is because, for small values of $p$, the POMDP-based and the LC policies transmit less frequently, while the baseline policies send updates more frequently. Actually, for a high-dynamic source, there is a high chance that the source will change its state in the next slot, and the currently transmitted status update cannot decrease the average cost at the next slot. For $p=0.6$, the energy-agnostic LC police’s performance becomes the same as the BO-RC policy; this is because, as shown in Fig. 6(a), the energy-agnostic LC policy starts to send updates for large $p$ whenever the battery has energy and the expected distortion cost is not zero, which is the main idea behind the BO-RC policy. For large values of $p$ (e.g., $p=0.7,p=0.8,p=0.9$), the performance of the BO-RC and energy-agnostic LC policies are very close to the POMDP-based and energy-aware LC policies and outperform the BO policy. This is because when the source state changes slowly, i.e., for large values of $p$, and the channel condition is good, most of the time the expected cost is zero, thus, these policies do not send updates frequently which saves energy for future transmissions. However, the BO policy sends updates independent of the cost value.
0.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.90.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9Source self transition probability, $p$Average cost, $\bar{C}$ policyBO-RC policyEnergy-agnostic LC policyEnergy-aware LC policyPOMDP-based policy
@@
Figure 7: The average cost $\bar{C}$ of different policies with respect to $p$ where $N=3$, $p_{s}=0.6$, $p_{f}=0.6$, $\mu=0.5$, and $B=3$.
@@
Fig. 8 shows the average cost $\bar{C}$ versus the energy arrival rate $\mu$ for different policies. As it can be seen, for small values of $\mu$, the POMDP-based policy outperforms other policies, while for large values of $\mu$, the performance of the policies except for the BO policy is very close. This is because in cases where the energy arrival rates $\mu$ are small, there are fewer transmission opportunities, and the other policies utilize these opportunities less effectively than the POMDP-based policy. Moreover, the energy-aware LC policy outperforms the energy-agnostic LC, BO-RC, and BO policies. This is because the energy-aware LC policy uses energy more effectively.
0.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.60.60.60.60.70.70.70.70.80.80.80.8Energy arrival rate, $\mu$Average cost, $\bar{C}$ policyBO-RC policyEnergy-agnostic LC policyEnergy-aware LC policyPOMDP-based policy
@@
Figure 8: The average cost $\bar{C}$ of different policies with respect to the energy arrival rate $\mu$, where $p=0.7$, $N=3$, $p_{f}=0.6$, $p_{s}=0.6$, and $B=3$.
@@
Fig. 9 demonstrates the average cost versus the number of source’s states $N$ under different policies. From the figure, the average cost increases with $N$, as expected. Moreover, the POMDP-based and energy-aware LC policies outperform other policies, while for large $N$ (e.g., $N=4$ and $N=5$), the POMDP-based policy outperforms the energy-aware LC policy. This is because increasing the number of states, in general, leads to a larger estimation error. Consequently, the system needs to use the available energy effectively by considering the state of the system in future time slots. However, except for the POMDP-based and energy-aware LC policies, the other policies are agnostic to the energy arrival rate.
22223333444455550.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.21.41.41.41.4Number of source states, $N$Average cost, $\bar{C}$ policyBO-RC policyEnergy-agnostic LC policyEnergy-aware LC policyPOMDP-based policy
@@
Figure 9: The average cost $\bar{C}$ of different policies with respect to numbers of source’s state $N$, where $p=0.7$, $N=3$, $p_{s}=0.6$, $p_{f}=0.2$, $\mu=0.2$, and $B=3$.
@@
In Fig. 10, we examine the impact of the probability of successful reception of ACK $p_{f}$ for different policies. As can be seen, the average cost decreases by an increase in $p_{f}$ except for the BO policy. This is because the number of slots that the sink is observable by the transmitter increases, which in turn, reduces the uncertainty in the system, leading to effective decision-making for those policies that consider the value of the belief. However, the BO policy makes decisions only based on the energy level of the battery, and thus, the reduction in the uncertainty of the system does not change its performance. In general, this figure demonstrates 1) the effectiveness of the proposed POMDP-based policy under poor feedback channel conditions and 2) its applicability for low-resource wireless networks, where having high-reliable communication is more costly.
0.050.050.050.050.250.250.250.250.50.50.50.50.750.750.750.7511110.630.630.630.630.650.650.650.650.670.670.670.670.70.70.70.70.720.720.720.72Probability of success in feedback channel, $p_{f}$verage cost, $\bar{C}$ policyBO-RC policyEnergy-agnostic LC policyEnergy-aware LC policyPOMDP-based policy
@@
Figure 10: The average cost $\bar{C}$ of different policies with respect to the probability of successful reception of ACK $p_{f}$, where $p=0.7$, $N=3$, $p_{s}=0.4$, $\mu=0.5$, and $B=3$.
@@
Finally, Fig. 11 depicts the average cost with respect to the battery’s capacity for different energy arrival rates. Here, without loss of generality, we use the POMDP-based policy. As expected, the average cost decreases by an increase in $B$. This is because a larger capacity of the battery allows the transmitter to store more energy, compensating for periods of inconsistent energy reception and ensuring continuous operation. The influence of the battery’s capacity on performance diminishes with increasing energy arrival rate. This is because, at a higher energy arrival rate, the transmitter receives energy more frequently. Therefore, the system can rely more on frequent energy arrival rather than large battery capacities.
11113333666699990.50.50.50.50.550.550.550.550.60.60.60.60.650.650.650.65The battery’s capacity, $B$Average cost, $\bar{C}$$\mu=0.3$$\mu=0.5$$\mu=0.7$
@@
Figure 11: The average cost $\bar{C}$ for different values of the battery’s energy arrival $\mu$ with respect to the battery’s capacity $B$, where $p=0.7$, $N=3$, $p_{s}=0.6$, and ${p_{f}=0.7}$.
@@
In Table I, we compare the computational complexity of the proposed policies in the offline phase, i.e., the initial processing time to find a policy, and 2) in the online phase, i.e., running time to find the optimal action at each slot. The computational complexity of the RVIA is at most $\mathcal{O}(|\mathcal{A}_{l}||\mathcal{L}|^{2})$ [8], where $|\mathcal{A}_{l}|$ is the size of action space and $|\mathcal{L}|$ is the size of belief-state space. The size of the action space is $2$. If we start from a reset belief, after one transmission attempt without receiving ACK, i.e., $m=1$, there are $N-1$ possible beliefs, and for $m=2$, there are $N$ possible beliefs. This procedure can start from any reset belief. Thus, the number of beliefs in the belief space is ${N(N-1)N^{m-1}=(N-1)N^{m}}$. Considering the other elements of the belief-state, the size of the belief-state space becomes $2B(N-1)N^{m+2}$. Therefore, the computational complexity for the offline phase of the POMDP-based policy is $\mathcal{O}([B(N-1)N^{m+2}]^{2})$. In the online phase, the policy needs to find the action from a state-action table, thus the computational complexity is ${\mathcal{O}(1)}$. The energy-agnostic LC policy does not have an offline phase. In the online phase, it needs to compare the cost associated with two actions, thus the computational complexity is ${\mathcal{O}(1)}$. According to the energy-aware LC policy, one needs to find the best $\gamma$ that provides the minimum average cost, and thus, the complexity for the offline phase is ${\mathcal{O}(1)}$. In the online phase, it is required to compare the cost associated with the two actions, and thus, the computational complexity is ${\mathcal{O}(1)}$.
@@
POMDP-based $\mathcal{O}([B(N-1)N^{m+2}]^{2})$ $\mathcal{O}(1)$
Energy-agnostic LC - $\mathcal{O}(1)$
Energy-aware LC $\mathcal{O}(1)$ $\mathcal{O}(1)$